# A weakly-modeled search for compact binary coalescences in Einstein Telescope

### v1

@@ Abstract
Einstein Telescope (ET) is a project of third generation gravitational-wave (GW) detector with a planned sensitivity $\sim 10$ times better than current detectors such as Advanced LIGO and Advanced Virgo. The high rate of GW signals expected in the data will pose several data analysis challenges, like the ability to disentangle overlapping signals or the need to dimension the computational resources required to treat all the candidate events. We explore the behaviour and the performances of a data analysis pipeline designed to search for unmodelled GW signals with duration $\sim 1-1000$ s on a mock dataset that consists of $1$ month of data following ET design sensitivity on top of which is added a realistic distribution of compact binary coalescence (CBC) signals. Unmodelled searches are intrinsically less sensitive to CBC signals than template-based searches, but are computationally cheaper and more robust to uncertainties in the waveforms. This search recovers $38\%$ of the total number of injected binary black hole (BBH) signals, including $89\%$ of the systems with a total mass above $100$ M, as well as the majority of binary neutron star (BNS) signals closer than $850$ Mpc ($z=0.17$). It is also able to estimate the chirp mass of the recovered BNS with an average precision of $1.3\%$. Therefore, we show that this unmodelled search is able to detect a substantial amount of CBC events at a relatively low computational cost, which makes it interesting for low-latency analyses and independent validation of detections made by matched filtering pipelines. We also find that the presence of many CBC signals only marginally impacts the sensitivity of the search to other kinds of unmodelled long-duration transient signals, by $\sim 3\%$ in average.
]compiled August 23, 2024
@@ I Introduction
Third-generation gravitational-wave (GW) detectors such as Einstein Telescope (ET) Punturo et al. (2010) and Cosmic Explorer Reitze et al. (2019); Evans et al. (2021) are expected to become operational in the mid-$2030s$. They aim to improve the sensitivity by an order of magnitude compared to the current generation of advanced detectors (Advanced LIGO Abbott et al. (2015), Advanced Virgo Acernese et al. (2015) and KAGRA Akutsu et al. (2019), commonly referred to as LVK), and to extend the sensitive frequency band at both low and high frequencies. In its current planed design, ET will be made of three nested detectors forming an equilateral triangle with $10$ km sides. Each detector will consist in two dual-recycled Fabry-Perot-Michelson interferometers tuned to be sensitive to high and low frequencies respectively, and the whole infrastructure will be built underground to limit the impact of seismic noise.
@@
The sheer amount of signals in the data, a fraction of which with signal-to-noise ratio (SNR) above $20$ Allen et al. (2012), will introduce new data analysis challenges. Because of the extended sensitivity at low frequency, CBC signals will be present for a longer time in the frequency band of the detectors, up to several hours or days in the case of BNS Regimbau and Hughes (2009). A fraction of these signals will overlap each other, which may affect the ability of data analysis pipelines to detect and to estimate the parameters of such signals Regimbau et al. (2012); Janquart et al. (2023); Johnson et al. (2024). The sum of all CBC events that are not individually resolvable will also create a confusion background that will complicate the observation of a stochastic GW background of cosmological origin, which is one of the big targets of GW astronomy Regimbau (2011); Regimbau et al. (2014); Caprini and Figueroa (2018); Christensen (2019); van Remortel et al. (2023) (see Abbott et al. (2021a) for the most recent constraints provided by LVK). Methods to subtract this background are therefore being developed Sachdev et al. (2020); Zhou et al. (2023); Zhong et al. (2023). Finally, the computational resources used to analyze the data will need to be scaled up to account for the larger bandwidth of the detector and a rate of events orders of magnitude higher than with current detectors Lenon et al. (2021); Bagnasco et al. (2023). In the context of multi-messenger astronomy, it is also crucial to provide rapid signal detections and estimations of the spatial position of the source, in order to trigger space- and Earth-based electromagnetic observations of the same source as soon as possible Abbott et al. (2017). Therefore, robust and rapid search algorithms will be needed to cope with the rate of detectable signals, such as the one proposed in Miller et al. (2024).
@@
In this work, we investigate whether an unmodeled search algorithm can be used to detect CBC signals in ET and to provide an estimation of the chirp mass, and we study the impact of the CBC foreground on the detectability of other types of unmodeled transients. We use PySTAMPAS Macquet et al. (2021), a data analysis pipeline designed to search for long-duration ($\sim 1-1000$ s) GW bursts, to analyse the latest ET mock data challenge (MDC), that consists of one month of data simulating ET noise and a realistic distribution of CBC events. The paper is organized as follows. In Sec. II, we present the dataset and describe the search algorithm we use to analyze it. Results of these analyzes are reported in Sec. III. In Sec. IV, we study the ability of the pipeline to detect non-CBC transient signals in the presence of a CBC foreground. We discuss the results obtained and the takeways of this study in Sec. V
@@ II.1 Description of the dataset
This MDC contains $31$ days of data simulating the output of the three V-shaped nested interferometers forming ET in the triangular configuration. For each detector, the data are the sum of colored Gaussian noise and the detector’s response to a distribution of GW signals from CBC sources. The dataset was generated with the with the MDC_Generation pipeline Regimbau (2024), an updated version of the code used for past MDC Regimbau et al. (2012, 2014); Meacher et al. (2016). To compute the response of the three detectors to GW signals, it is assumed that they are located at the current EGO site near Cascina in Italy. The PSD of the noise is the reference one defined in Branchesi et al. (2023) for $10$ km arms in the xylophone configuration (combining high and low-frequency interferometers). The noise is purely Gaussian, colored with the PSD, and uncorrelated between the detectors.
@@ II.2 Search methodology
We analyze this dataset with PySTAMPAS, a data analysis pipeline designed to search for un-modelled GW transients with duration ranging from few seconds to minutes. A detailed overview of the pipeline and its implementation can be found in Macquet et al. (2021). The following describes its basic workflow and the configuration used for this analysis.
@@
The first one is a seed-based clustering algorithm, which identifies groups of bright pixels by time-frequency proximity to form clusters. Seed-based algorithms are sensitive to any signal morphology, provided that it is contiguous in the tf-map and loud enough to generate high SNR pixels. They are therefore suited for generic unmodelled searches with minimal assumptions on the morphology of the signals searched. We use the burstegard algorithm Prestegard (2016), that is commonly used for long-duration burst searches in LVK data Abbott et al. (2019, 2021b).
@@ III.1 Background estimation
To assess the significance of candidate events found by the search, one needs to estimate the rate of background triggers, i.e triggers that are generated by noise fluctuations in the detectors. In general, GW detectors’ noise is not Gaussian and presents many transient features that generate noise triggers whose distribution cannot be predicted analytically. Therefore, the background distribution must be estimated empirically. As many other GW search algorithms, PySTAMPAS implements the time-slides method, which consists in shifting the data between the detectors by an amount of time greater than the maximal duration of the GW signal searched. Therefore, the properties of the noise in individual detectors are preserved, while any GW signal present in the data appears as incoherent. This method also allows to simulate a large amount of background data by replicating the experiment with different values of the time shift. The cumulative rate of events found in time-shifted data is used as an estimator of the false-alarm rate (FAR) of the search.
@@
In PySTAMPAS, clusters of pixels extracted from each detector are time-shifted with respect to the other detectors’ data to estimate the background. For this search, we use $1216$ different values for the time shift, simulating $\sim 102$ years of background noise. The cumulative rate of background triggers as a function of their ranking statistic $p_{\Lambda}$ is represented by the blue curve in Fig. 1. It is significantly higher than the rate we would expect for pure Gaussian noise, which is shown by the orange curve, and even higher than what we typically observe in realistic LVK searches, where the noise is dominated by loud transient artifacts Macquet et al. (2021). Due to the high rate of high SNR signals in the dataset, most of the clusters extracted from single-detector tf-maps actually correspond to GW signals, and there is a significant probability that a signal from one detector is time-shifted on top of another signal in the other detectors, creating coherent pixels that increase the overall significance of the trigger. The main consequence is an apparent reduction of the sensitivity of the search, as the significance of events found in coincidence is reduced.
@@
This result illustrates the (well-known) limits of the time slides method when the data contains many loud signals, as it is expected for third-generation GW detectors such as ET. However, the triangular configuration considered here for ET offers a convenient way to bypass this issue in the form of the null stream. In general, the null stream is a linear combination of the output of the detectors that cancels a given GW signal. Such combination always exists for an array of three or more detectors, as was first demonstrated by the authors of Gürsel and Tinto (1989). In the general case of arbitrarily located and oriented detectors, that combination depends on the sky position of the source, which must therefore be known a priori (see e.g Chatterji et al. (2006) or Sutton et al. (2010) for the mathematical details on the computation of the null stream). However, in the particular case of a triangular configuration, the sum of the responses of the three detectors to any GW signal is always zero Freise et al. (2009); Regimbau et al. (2012). This allows one to build a single null stream that effectively cancels all GW signals present in the data, assuming the three interferometers are perfectly calibrated.
@@ III.2 Search for generic transients with burstegard
In the end, $2620$ out of $3035$ triggers ($85\%$) are uniquely matched with one injection. Less than $1\%$ of the total number of significant triggers are not matched with any injection. This can happen for two reasons: first, because the analysis windows are only $512$ s long, a signal may be split between two or more windows and generate triggers in several windows. Because the matching procedure looks only at the coalescence time, triggers that correspond to the earlier parts of the signal will not be matched. Also, burstegard may cluster noise pixels at the end of a signal, leading to the end time of the trigger being overestimated, sometimes by more than the threshold of $20$ s chosen for matching. We consider these reasons as intrinsic limitations of the seed-based clustering algorithm, and therefore we decide to discard these triggers even if they correspond to actual GW signals. These limitations are illustrated in Fig. 2, which shows examples of triggers extracted by PySTAMPAS: in both cases, two high-SNR signals have been reconstructed as a single event because they overlap in time and frequency. In the first case, because the merger time of the BNS is not located in the same analysis window, none of these two injections have been associated with the trigger. Conversely, one injected signal can be matched with several triggers. Because of the seed-based nature of the clustering algorithm, a signal may be split into several clusters. If this split happens close to the merger time, the two triggers will both have end times close to the reference coalescence time and will then be associated with the same injection. We note that in general in MDCs, the procedure to associate recovered triggers to injected signals must rely on at least some arbitrary choices. For this particular study, we have chosen to use the difference between the trigger end time and the coalescence time of the CBC as the main criterion, because the seed-based clustering algorithm tends to better recover the merger part of the CBC. We find that a tolerance window of ±$\pm 20$ s is sufficiently short to suppress the risk of false association while minimizing the number of injections missed.
@@
Among the $2620$ injections recovered, $2567$ are labelled as BBH, $44$ as BNS, and $8$ as NSBH. The farthest event recovered is at a redshift $z=13.1$. It is a BBH with a total mass of $62.8$ M in the source frame ($885$ M in the detector frame). The distribution of recovered and missed signals in mass and redshift is shown in Fig. 3 for the two clustering algorithms used. In total, this search recovers $38\%$ of the total number of injected BBH, including $65\%$ of the ones with a total mass in the source frame above $50$ M, and $89\%$ above $100$ M. High-mass BBH have a relatively short duration (a few seconds or less) in the frequency band of the detectors, so they are spread among a smaller number of time-frequency pixels than lower mass CBC, which are also less energetic. Therefore, they generate high SNR pixels that are picked up by the clustering algorithm, which explains why they are well recovered by PySTAMPAS, despite the pipeline being tuned to detect longer duration signals. It is therefore likely that other unmodelled search algorithms designed to target signals with duration ranging from milliseconds to few seconds would be able to recover an even larger fraction of BBH.
Figure 3: Redshift of injected CBC signals as a function of their total mass in the source frame. Green dots represent signals recovered using the seed-based burstegard clustering algorithm, while blue crosses are signals recovered by the seedless algorithm. Red crosses represent missed injections. The blue curve represents the maximum redshift at which an equal-mass, non-spinning, quasicircular binary coalescence would be detectable with an optimal SNR $\geq 8$.
@@ III.3 Search for chirp-like signals
For each tf-map, we test $100$ values of $\mathcal{M}_{c}$ logarithmically spaced between $1$ and $5$ M, and $1024$ values of $t_{c}$ separated by $0.5$ s. We do not test values of $\mathcal{M}_{c}$ above $5$ M as we are mostly interested in improving the recovery of BNS signals for which the seed-based clustering algorithm is not well suited. For each template, the corresponding cluster is the set of pixels $\Gamma$ that overlap the graph $(t,f(t))$. We define the (single-detector) ranking statistic
@@
In order to test the algorithm in a controlled way, we estimate the detection efficiency of the search for a reference $1.4-1.4$ M non-spinning BNS on the noise used for this MDC and compare the two clustering algorithms. The detection efficiency as a function of the injected distance is shown in Fig. 4 for a FAR of $1$ per year. The horizon, defined as the distance for which the detection efficiency reaches $50\%$, is $382$ Mpc for burstegard, and goes up to $717$ Mpc when using this implementation of seedless clustering, an increase of $88\%$. The distance at which $50\%$ of events have an SNR $\geq 8$ in at least one detector is $1275$ Mpc.
Figure 4: Detection efficiency as a function of the distance for a $1.4-1.4$ M BNS (IMRPhenomPv2 approximant) with seed-based (orange) and seedless (blue) clustering. The green curve represents the fraction of signals for which the matched filter SNR is above $8$ in at least one detector.
On the MDC data, we recover $81$ BNS in total, $84\%$ more than with burstegard. The farthest BNS recovered is at redshift $z=0.68$, and $50\%$ of BNS below $z=0.17$ are recovered. The distribution of BNS recovered using the seedless clustering algorithm is shown by the blue crosses in Fig. 3.
Finally, we compare the chirp mass recovered by the algorithm to the values injected for each of the $81$ BNS recovered (Fig. 5). The root mean square error (RMS) between the two values is $0.04$ M, which corresponds to an average relative error of $1.3\%$. We note that the chirp mass estimated here is the one in the detector frame. To estimate the mass in the source frame, an independent estimation of the redshift would be needed, which is outside of the scope of this method and would probably have to rely on full Bayesian inference with CBC waveform models triggered by the PySTAMPAS candidate Ashton et al. (2019).
The RMS for the coalescence time is $0.39$ s, which is consistent with the $0.5$ s separation between the different values of $t_{c}$ tested.
Figure 5: Relative error on the reconstructed chirp mass in the detector frame as a function of the injected one for the $81$ BNS recovered by the search using seedless clustering.

### v2

@@ Abstract
Einstein Telescope (ET) is a third-generation gravitational-wave (GW) detector project which aims to reach a sensitivity $\sim 10$ times larger than current GW detectors such as Advanced LIGO and Advanced Virgo. The high rate of GW signals expected in the data will pose several data analysis challenges, such as the ability to disentangle overlapping signals, or the computational resources required to treat all the candidate events. We explore the behaviour and the performances of PySTAMPAS, a data analysis pipeline designed to search for unmodeled GW signals with duration $\sim 1-1000$ s on a mock dataset that consists of $1$ month of data following ET design sensitivity on top of which is added a realistic distribution of compact binary coalescence (CBC) signals. Unmodeled searches are intrinsically less sensitive to CBC signals than template-based searches, but are computationally cheaper and more robust to uncertainties in the waveforms. This search recovers $38\%$ of the total number of injected binary black hole (BBH) signals, including $89\%$ of the systems with a total mass above $100$ M, as well as the majority of binary neutron star (BNS) signals closer than $850$ Mpc ($z=0.17$). It is also able to estimate the chirp mass of the recovered BNS with an average precision of $1.3\%$. Therefore, we show that this unmodeled search is able to detect a substantial amount of CBC events at a relatively low computational cost, which makes it interesting for low-latency analyses and independent validation of detections made by matched filtering pipelines. We also find that the presence of many CBC signals only marginally impacts the sensitivity of the search to other kinds of unmodeled long-duration transient signals, by $\sim 3\%$ in average.
]compiled December 19, 2024
@@ I Introduction
Third-generation gravitational-wave (GW) detectors such as Einstein Telescope (ET) Punturo et al. (2010) and Cosmic Explorer Reitze et al. (2019); Evans et al. (2021) are expected to become operational in the mid-$2030s$. They aim to improve the sensitivity by an order of magnitude compared to the current generation of advanced detectors (Advanced LIGO Abbott et al. (2015), Advanced Virgo Acernese et al. (2015) and KAGRA Akutsu et al. (2019), commonly referred to as LVK), and to extend the sensitive frequency band at both low and high frequencies. In its current planed design, ET will be made of three nested detectors forming an equilateral triangle with $10$ km sides. Each detector will consist of two dual-recycled Fabry-Perot-Michelson interferometers tuned to be sensitive to high and low frequencies respectively, and the whole infrastructure will be built underground to limit the impact of seismic noise.
@@
The sheer number of signals in the data, a fraction of which with signal-to-noise ratio (SNR) above $20$ Allen et al. (2012), will introduce new data analysis challenges. Because of the extended sensitivity at low frequency, CBC signals will be present for a longer time in the frequency band of the detectors, up to several hours or days in the case of BNS Regimbau and Hughes (2009). A fraction of these signals will overlap each other, which may affect the ability of data analysis pipelines to detect and to estimate the parameters of such signals Regimbau et al. (2012); Janquart et al. (2023); Johnson et al. (2024); Pizzati et al. (2022); Relton and Raymond (2021); Samajdar et al. (2021); Antonelli et al. (2021); Relton et al. (2022). The sum of all CBC events that are not individually resolvable will also create a confusion background that will complicate the observation of a stochastic GW background of cosmological origin, which is one of the big targets of GW astronomy Regimbau (2011); Regimbau et al. (2014); Caprini and Figueroa (2018); Christensen (2019); van Remortel et al. (2023) (see Abbott et al. (2021a) for the most recent constraints provided by LVK). Methods to subtract this background are therefore being developed Sachdev et al. (2020); Zhou et al. (2023); Zhong et al. (2023). Finally, the computational resources used to analyze the data will need to be scaled up to account for the larger bandwidth of the detector and a rate of events orders of magnitude higher than with current detectors Lenon et al. (2021); Bagnasco et al. (2023). In the context of multi-messenger astronomy, it is also crucial to provide rapid signal detections and estimations of the spatial position of the source, in order to trigger space- and Earth-based electromagnetic observations of the same source as soon as possible Abbott et al. (2017). Therefore, robust and rapid search algorithms will be needed to cope with the rate of detectable signals, such as the one proposed in Miller et al. (2024).
@@
PySTAMPAS Macquet et al. (2021) is a data analysis pipeline originally designed to search for long-duration ($\sim 1-1000$ s) GW bursts in LVK data. In this work, we investigate whether this specific search algorithm could also be used to detect CBC signals in ET and to provide an estimation of the chirp mass, and we study the impact of the CBC foreground on the detectability of other types of unmodeled transients. The choice of a search algorithm that targets signals with duration $>1$ s is motivated by the extended duration of CBC signals due to the improved sensitivity of ET at low frequency. This work takes place in the context of the latest ET mock data challenge (MDC) which aims to provide a common dataset to investigate how current data analysis techniques would perform in the context of ET, identify the science cases, and test new data analysis and parameter estimation methods. In this paper, we analyse this MDC that consists of one month of data simulating ET noise and a realistic distribution of CBC events, with PySTAMPAS. The paper is organized as follows. In Sec. II, we present the dataset and describe the search algorithm we use to analyze it. Results of these analyzes are reported in Sec. III. In Sec. IV, we study the ability of the pipeline to detect non-CBC transient signals in the presence of a CBC foreground. We discuss the results obtained and the takeways of this study in Sec. V
@@ II.1 Description of the dataset
This MDC contains $31$ days of data simulating the output of the three V-shaped nested interferometers forming ET in the triangular configuration. For each detector, the data are the sum of colored Gaussian noise and the detector’s response to a distribution of GW signals from CBC sources. The dataset was generated with the with the MDC_Generation pipeline Regimbau (2024), a code already used for past MDC Regimbau et al. (2012, 2014); Meacher et al. (2016). To compute the response of the three detectors to GW signals, it is assumed that they are located at the current EGO site near Cascina in Italy. The PSD of the noise is the reference one defined in Branchesi et al. (2023) for $10$ km arms in the xylophone configuration (combining high and low-frequency interferometers). The noise is purely Gaussian, colored with the PSD, and uncorrelated between the detectors.
@@ II.2 Search methodology
We analyze this dataset with PySTAMPAS, a data analysis pipeline designed to search for un-modeled GW transients with duration ranging from few seconds to minutes. A detailed overview of the pipeline and its implementation can be found in Macquet et al. (2021). The following describes its basic workflow and the configuration used for this analysis.
@@
The first one is a seed-based clustering algorithm, which identifies groups of bright pixels by time-frequency proximity to form clusters. Seed-based algorithms are sensitive to any signal morphology, provided that it is contiguous in the tf-map and loud enough to generate high SNR pixels. They are therefore suited for generic unmodeled searches with minimal assumptions on the morphology of the signals searched. We use the burstegard algorithm Prestegard (2016), that is commonly used for long-duration burst searches in LVK data Abbott et al. (2019, 2021b).
@@ III.1 Background estimation
To assess the significance of candidate events found by the search, one needs to estimate the rate of background triggers, i.e., triggers that are generated by noise fluctuations in the detectors. In general, GW detectors’ noise is not Gaussian and presents many transient features that generate noise triggers whose distribution cannot be predicted analytically. Therefore, the background distribution must be estimated empirically. As many other GW search algorithms, PySTAMPAS implements the time-slides method, which consists in shifting the data between the detectors by an amount of time greater than the maximal duration of the GW signal searched. Therefore, the properties of the noise in individual detectors are preserved, while any GW signal present in the data appears as incoherent. This method also allows to simulate a large amount of background data by replicating the experiment with different values of the time shift. The cumulative rate of events found in time-shifted data is used as an estimator of the false-alarm rate (FAR) of the search.
@@
In PySTAMPAS, clusters of pixels extracted from each detector are time-shifted with respect to the other detectors’ data to estimate the background. For this search, we use $1216$ different values for the time shift, simulating $\sim 102$ years of background noise. The cumulative rate of background triggers as a function of their ranking statistic $p_{\Lambda}$ is represented by the blue curve in Fig. 1 for the E1-E2 baseline. It is significantly higher than the rate we would expect for pure Gaussian noise, which is shown by the orange curve, and even higher than what we typically observe in realistic LVK searches, where the noise is dominated by loud transient artifacts Macquet et al. (2021). Due to the high rate of high SNR signals in the dataset, most of the clusters extracted from single-detector tf-maps actually correspond to GW signals, and there is a significant probability that a signal from one detector is time-shifted on top of another signal in the other detectors, creating coherent pixels that increase the overall significance of the trigger. The main consequence is an apparent reduction of the sensitivity of the search, as the significance of events found in coincidence is reduced.
@@
This result illustrates the (well-known) limits of the time slides method when the data contains many loud signals, as it is expected for third-generation GW detectors such as ET. However, the triangular configuration considered here for ET offers a convenient way to bypass this issue in the form of the null stream. In general, the null stream is a linear combination of the output of the detectors that cancels a given GW signal. Such combination always exists for an array of three or more detectors, as was first demonstrated by the authors of Gürsel and Tinto (1989). In the general case of arbitrarily located and oriented detectors, that combination depends on the sky position of the source, which must therefore be known a priori (see e.g Chatterji et al. (2006) or Sutton et al. (2010) for the mathematical details on the computation of the null stream). However, in the particular case of a triangular configuration, the sum of the responses of the three detectors to any GW signal is always zero Freise et al. (2009); Regimbau et al. (2012). This allows one to build a single null stream that effectively cancels all GW signals present in the data, assuming the three interferometers are perfectly calibrated Schutz and Sathyaprakash (2020).
@@ III.2 Search for generic transients with burstegard
In the end, $2620$ out of $3035$ triggers ($85\%$) are uniquely matched with one injection. Less than $1\%$ of the total number of significant triggers are not matched with any injection. This can happen for two reasons: first, because the analysis windows are only $512$ s long, a signal may be split between two or more windows and generate triggers in several windows. Because the matching procedure looks only at the coalescence time, triggers that correspond to the earlier parts of the signal will not be matched. In theory, it would be possible to increase the duration of analysis windows to be able to better reconstruct hours-long BNS signals, but that would be at the cost of an increased computing time, as the computational cost of the clustering algorithms used in this search increases with the number of pixels in the time-frequency map. Also, burstegard may cluster noise pixels at the end of a signal, leading to the end time of the trigger being overestimated, sometimes by more than the threshold of $20$ s chosen for matching. We consider these reasons as intrinsic limitations of the seed-based clustering algorithm, and therefore we decide to discard these triggers even if they correspond to actual GW signals. These limitations are illustrated in Fig. 2, which shows examples of triggers extracted by PySTAMPAS: in both cases, two high-SNR signals have been reconstructed as a single event because they overlap in time and frequency. In the first case, because the merger time of the BNS is not located in the same analysis window, none of these two injections have been associated with the trigger. Conversely, one injected signal can be matched with several triggers. Because of the seed-based nature of the clustering algorithm, a signal may be split into several clusters. If this split happens close to the merger time, the two triggers will both have end times close to the reference coalescence time and will then be associated with the same injection. We note that in general in MDCs, the procedure to associate recovered triggers to injected signals must rely on at least some arbitrary choices. For this particular study, we have chosen to use the difference between the trigger end time and the coalescence time of the CBC as the main criterion, because the seed-based clustering algorithm tends to better recover the merger part of the CBC. We find that a tolerance window of ±$\pm 20$ s is sufficiently short to suppress the risk of false association while minimizing the number of injections missed.
@@
Among the $2620$ injections recovered, $2567$ are labelled as BBH, $44$ as BNS, and $8$ as NSBH. The farthest event recovered is at a redshift $z=13.1$. It is a BBH with a total mass of $62.8$ M in the source frame ($885$ M in the detector frame). The distribution of recovered and missed signals in mass and redshift is shown in Fig. 3 for the two clustering algorithms used. In total, this search recovers $38\%$ of the total number of injected BBH, including $65\%$ of the ones with a total mass in the source frame above $50$ M, and $89\%$ above $100$ M. High-mass BBH have a relatively short duration (of the order of a few seconds) in the frequency band of the detectors, so they are spread among a smaller number of time-frequency pixels than lower mass CBC, which are also less energetic. Therefore, they generate high SNR pixels that are picked up by the clustering algorithm, which explains why they are well recovered by PySTAMPAS, despite the pipeline being tuned to detect longer duration signals. It is therefore likely that other unmodeled search algorithms designed to target signals with duration ranging from milliseconds to few seconds would be able to recover an even larger fraction of BBH. In LVK data, these high mass BBH have even shorter durations (less than $1$ s), and are therefore not detectable at all by the pipeline.
When using the time-slides method to estimate the background instead of the null stream, the threshold on the ranking statistic $p_{\Lambda}$ corresponding to a FAR of $1$ per year increases, as shown by the blue curve in Fig. 1, which in turn reduces the number of signals effectively recovered from $2620$ to $1411$, a decrease of $46\%$ for the BBH signals.
Figure 3: Redshift of injected CBC signals as a function of their total mass in the source frame. Green dots represent signals recovered using the seed-based burstegard clustering algorithm, while blue crosses are signals recovered by the seedless algorithm. Red crosses represent missed injections. The blue and orange curves representr the maximum redshift at which an equal-mass, non-spinning, quasicircular binary coalescence would be detectable with an optimal SNR $\geq 8$ in ET and Advanced LIGO at design sensitivity respectively.
@@ III.3 Search for chirp-like signals
For each tf-map, we test $100$ values of $\mathcal{M}_{c}$ logarithmically spaced between $1$ and $5$ M, and $1024$ values of $t_{c}$ separated by $0.5$ s. We also run the same search with $100$ values of $\mathcal{M}_{c}$ spaced between $5$ and $200$ M to determine whether this algorithm is also sensitive to BBH and NSBH. For each template, the corresponding cluster is the set of pixels $\Gamma$ that overlap the graph $(t,f(t))$. We define the (single-detector) ranking statistic
@@
In order to test the algorithm in a controlled way, we estimate the detection efficiency of the search for a reference $1.4-1.4$ M non-spinning BNS on the noise used for this MDC and compare the two clustering algorithms. The detection efficiency as a function of the injected distance is shown in Fig. 4 for a FAR of $1$ per year. The horizon, defined as the distance for which the detection efficiency reaches $50\%$, is $382$ Mpc for burstegard, and goes up to $717$ Mpc when using this implementation of seedless clustering, an increase of $88\%$. The distance at which $50\%$ of events have an SNR $\geq 8$ in at least one detector is $1275$ Mpc. As a comparison, we do the same study on Gaussian noise following Advanced LIGO design sensitivity and find a horizon distance of $53$ Mpc using burstegard, a decrease by a factor $\sim 7$ which is consistent with the relative sensitivities of the detectors.
Figure 4: Detection efficiency as a function of the distance for a $1.4-1.4$ M BNS (IMRPhenomPv2 approximant) with seed-based (orange) and seedless (blue) clustering. The green curve represents the fraction of signals for which the matched filter SNR is above $8$ in at least one detector. The red curve represents the detection efficiency for the same waveform on Gaussian noise following Advanced LIGO design sensitivity for comparison (only the .
On the MDC data, we recover $81$ BNS in total, $84\%$ more than with burstegard. The farthest BNS recovered is at redshift $z=0.68$, and $50\%$ of BNS below $z=0.17$ are recovered. However, we recover only $712$ BBH and NSBH, $72\%$ less than burstegard, which is expected because seed-based clustering performs better on shorter signals. The distribution of CBC recovered using the seedless clustering algorithm is shown by the blue crosses in Fig. 3. When the background is estimated with the time-slides method, only $15$ BNS are recovered, a decrease of $81\%$. As explained above, BNS typically generate fainter pixels than BBH in time-frequency maps, so the associated triggers have a lower ranking statistic, leading to more of them being discarded when the recovery threshold increases.
Finally, we compare the chirp mass and coalescence time recovered by the algorithm to the values injected for each recovered signal in Fig. 5. Given that the algorithm performs poorly for BBH and NSBH, we focus here on the search targeting chirp masses between $1$ and $5$ M. The root mean square error (RMS) between the two values is $0.04$ M for BNS, which corresponds to an average relative error of $1.3\%$. We note that the chirp mass estimated here is the one in the detector frame. To estimate the mass in the source frame, an independent estimation of the redshift would be needed, which is outside of the scope of this method and would probably have to rely on full Bayesian inference with CBC waveform models triggered by the PySTAMPAS candidate Ashton et al. (2019). The RMS for the coalescence time is $0.39$ s, which is consistent with the $0.5$ s separation between the different values of $t_{c}$ tested, but does not match the precision of $\sim 0.1$ s reached by matched-filtering pipelines.
Figure 5: Relative error on the reconstructed chirp mass in the detector frame (top) and on the reconstructed coalescence time (bottom) as a function of the injected one for the $81$ BNS signals recovered by the search using seedless clustering.